A 112 kg stone boulder rolls down a 67.8 m long ramp with an angle of inclination of 27.7° in 10.3 s. What is the power of gravity on the boulder?


1
Expert's answer
2021-03-08T08:23:52-0500

"P=\frac At=\frac{Fh}{t}=\frac{Fl\text{sin}\alpha}{t}=\frac{mgl\text{sin}\alpha}{t},"

"P=\frac{112\cdot 9.8\cdot 67.8\cdot 0.46}{10.3}\approx 3.3~\text{kW}."